ah42632:

An angle bisector of a triangle divides the opposite side of the triangle into segments 6 cm and 5 cm long. A second side of the triangle is 6.9 cm long. Find the longest and shortest possible lengths of the third side of the triangle. Round answers to the nearest tenth of a centimeter. A. 41.4 cm, 8.3 cm B. 30 cm, 5.8 cm C. 41.4 cm, 4.3 cm D. 8.3 cm, 5.8 cm

sophia333:

BD/DC=6/5=AB/AC, in which AC=6.9, then we solve for AB=AC*6/5=8.28

sophia333:

BD/DC=6/5=AB/AC, where AB=6.9, then we solve AC=AB*5/6=5.75,

sophia333:

So have you gotten the answers-

ah42632:

nope. still confused. srry

sophia333:

So what did i say the answer was to both those problems we just solved-

ah42632:

ohhhhhhh. I get it now. if i round up then it would be D. 8.3, 5.8

YECR:

BD/DC=6/5=AB/AC, in which AC=6.9, then we solve for AB=AC*6/5=8.28 like he said If you ever neeed help just ask one of us or tag anyone